OpenStudy (anonymous):

MEEDAL A rectangle has a length of 6 cm by 8 cm. Which is the length of a diagonal of the rectangle?

OpenStudy (anonymous):

@MotherOfGod

OpenStudy (anonymous):

its 10 correct?

OpenStudy (anonymous):

The diagonal of a rectangle splits the rectangle into two right triangles. The diagonal of the rectangle is the hypoteneuse of each one of the right triangles. The triangles have sides with 6 and 8cm. So you can apply the Pythagorean Theorem: \[x^2 = 6^2 + 8^2\]

OpenStudy (anonymous):

So, yes! Your answer is correct! x^2=100, so x=10!